Can someone help me on this equation?
I need to complete the order pair so that each is a solution for the given equation:
3x + 4y = 12
pair 1 is (8/3,)
pair 2 is (0,)
Thanks,
Sher

Answer by rapaljer(4671)

3x + 4y = 12
If x=8/3, then 3*8/3+4y = 12
8+4y=12
4y=4
y=1
(8/3,1)

If x=0, then 3*0+4y=12
4y=12
y=3
(0,3)

R^2 at SCC